# A Neutron Sensitive Detector Using 3D-Printed Scintillators

### v1

@@ 6 Conclusions
This work demonstrates the successful fabrication of a neutron-sensitive scintillating 3D-printed detector via the FDM method, using commercially accessible materials and equipment. The developed scintillator, based on established compounds like Li-6, offers a scalable and rapid production pathway for scintillators. Integration with a fast optical detector system confirmed sensitivity to both ionizing radiation and thermal neutrons, achieving a measured light yield of 30 photons/MeV. Furthermore, a novel neutron/gamma discrimination algorithm was implemented, enabling effective separation of neutron signals from gamma-ray backgrounds, validating the detector’s performance in mixed radiation fields. To improve the performance of the scintillator, we are exploring the use of perovskite materials. This will improve the light output of the scintillator, while preserving the low decay times.

### v2

@@ 6 Conclusions
This work demonstrates the successful fabrication of a neutron-sensitive scintillating 3D-printed detector via the FDM method, using commercially accessible materials and equipment. The developed scintillator, based on established compounds like Li-6, offers a scalable and rapid production pathway for scintillators. In combination with large commercially available 3D printers, this offers the opportunity to produce scintillators for a wide variety of uses, from small scintillators for handheld detectors to metre-scale scintillators for purposes such as neutrino detectors [22]. Integration with a fast optical detector system confirmed sensitivity to both ionizing radiation and thermal neutrons, achieving a measured light yield of 30 5 photons/MeV. Furthermore, a novel neutron/gamma discrimination algorithm was implemented, enabling effective separation of neutron signals from gamma-ray backgrounds, validating the detector’s performance in mixed radiation fields. To improve the performance of the scintillator, we are exploring the use of perovskite materials. This will improve the light output of the scintillator, while preserving the low decay times.